Circular No. 8305

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 DO_{29} (SPACEWATCH-LINEAR) An apparently asteroidal object found by both the Spacewatch and

LINEAR projects on CCD images taken on Feb. 17 (discovery observations below from MPEC 2004-D45 and MPS 10267) has been found to have a 5′′ centrally condensed coma (of mag R = 18.7) and a 15′′ tail in p.a. 220^{o}- 270^{o} on CCD images taken on Mar. 16.3–16.6 UT by S. S. Sheppard, Y. R. Fernandez, and D. Jewitt with the University of Hawaii 2.2-m reflector.

2004 UT  α_{2000}  δ_{2000}  Mag. Observer
Feb. 17.21711 10^{h}47^{m}22.^{s}60 +10^{o}49′35′′.4 19.8 Spacewatch
17.27872 10 47 20.87 +10 49 56.9 19.4 LINEAR

Additional astrometry, the following orbital elements, and an ephemeris appear on MPEC 2004-F09.

T = 2004 Oct. 7.5747 TT  ω = 41.^{o}3922
e = 0.451320  Ω = 147.8274 }2000.0
q = 4.091407 AU  i = 14.4752
a = 7.456814 AU n^{o} = 0.0484032 P = 20.4 years

SUPERNOVA 2004au IN MCG +04-42-2 Further to IAUC 8302, W. Li et al. report the LOSS discovery of

a supernova on unfiltered KAIT images taken on Mar. 9.55 (mag 18.8) and 11.55 UT (mag 18.7). SN 2004au is located at α = 17^{h}42^{m}02^{s}.98, δ = +25^{o}37′11′′.9 (equinox 2000.0), which is 4′′.2 east and 0′′.6 south of the center of MCG +04-42-2. Nothing was visible at this location on a KAIT image taken on 2003 Sept. 13.21 (limiting mag 19.5).

A. V. Filippenko, R. J. Foley, and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra obtained on Mar. 16 UT with the Keck I 10-m telescope (+ LRIS; range 310–920 nm) shows that the object is of type II, roughly 1 month past maximum brightness, with well-developed P-Cyg profiles. The recession velocity of the host galaxy, measured from Hα in superposed H II regions, is 7300 km/s. The ejecta expansion velocity, derived from the minimum of the Hα absorption line, is ∼ 6600 km/s.

2004 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green